OpenStudy (anonymous):

120 pounds of cement and 4 gallons of water, what is the ratio of cement to water?

OpenStudy (anonymous):

find how heavy 4 gallons of water is using its density

hero (hero):

If c = pounds of cement w = gallons of water: then the ratio of cement to water is \[\frac{c}{w} = \frac{120}{4} = \frac{(4)(30)}{4} = \frac{4}{4} \times \frac{30}{1}= \frac{30}{1}\]

OpenStudy (anonymous):

thanks hero!!!